OpenStudy (anonymous):

I have done the rest of the problem, but I need help on the last part.. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

Part 1 - f(x) = 2x - 4 g(x) = x + 4 / 2 Part 2 - f(x) = 2x - 4 Change f(x) for y y = 2x -4 Switch x and y x = 2y -4 Isolate y by adding 4 to both sides x + 4 = 2y Divide both sides by 2 x + 4 = 2y / 2 y = x + 4 / 2 Part 3 - g(f(x)) = (2x - 4) +4 / 2

OpenStudy (gorv):

well in each graph take x=0 and find f(0)

OpenStudy (gorv):

this will be one point (0,f(0))

OpenStudy (triciaal):

you did not post the original but it seem like you just have functions on the same graph. maybe 3 or 4 lines y = x is also a line

OpenStudy (gorv):

now take f(x)=0 and found x =?? this will be other point (x,0) join these 2 points

OpenStudy (anonymous):

I just find the graph hard to do.. Lol but I also don't know how to create a graph on the computer..

OpenStudy (gorv):

that will be our function

OpenStudy (anonymous):

so it would be like (2,0) and (4,0) ?

OpenStudy (anonymous):

What do you mean by the original? @triciaal

OpenStudy (gorv):

f(x)=2x-4 in this x=0 f(0)=-4 (0,-4)

OpenStudy (triciaal):

what was the question before you did anything?

OpenStudy (anonymous):

okay that kinda makes sense.. I think lol

OpenStudy (gorv):

second take f(x)=0 0=2x-4 2x=4 x=2 (2,0)

OpenStudy (gorv):

now we have 2 point but we have to find 5 point

OpenStudy (anonymous):

Well, the work that I added below my question was what leads up to making this graph.. One second

OpenStudy (gorv):

so similarly take small values for easy calculation

OpenStudy (gorv):

like x=1 find f(1)

OpenStudy (anonymous):

Okay, I think I had a guess on what the points would be. But, how am I going to create a graph?

OpenStudy (gorv):

like (0,2 )

OpenStudy (anonymous):

I do understand what you are saying, but I am not a graph person xD

OpenStudy (gorv):

took x=0 and y=2 find where they will meet

OpenStudy (anonymous):

I need to be able to attach a picture of the graph to the assignment but I have no clue how to create a graph on a computer..

OpenStudy (gorv):

on computer i dont have any idea

OpenStudy (gorv):

by hand i can help

OpenStudy (gorv):

different process for different software

OpenStudy (gorv):

use any software on net

OpenStudy (anonymous):

Crap /: Do you know how I would draw this graph out?

OpenStudy (anonymous):

I will draw yours out that you attach and take a picture of it, if you can do that?

OpenStudy (gorv):

like (2,2) and (5,4)

OpenStudy (gorv):

locate these point

OpenStudy (anonymous):

Okay, I see you found a site>?